I need help with this question of Differential Equation.

Thanks

Find and classify all singular points of the following differential equations. a) (x + 2)2(x - 1)y + 3(x - 1)y – 2(x + 2)y
Answer #1

ca(x) ?(2+) yllt 340–1341- 24*2)y=0 y + 34 2 ED (242)2 (2+2) (24) P = 3 2 cata) 2, QE (2+2) (2-1) 2= -2 & 2= I because P E Q